OpenStudy (anonymous):

What is the simplified form of the expression 5(14-2)2 over 2 A 60 B 30 C 72 D 360

OpenStudy (goldphenoix):

Think of PEMDAS. According to PEMDAS, you perform parenthesis. (14-2) = 12. [5(12)2]/2. Next multiply from left to right. 5*12*2 = 120 Finally divide. 120/2 = 60